OpenStudy (anonymous):

multiply 4 2/3 x 3 and 5/6 can someone please explain how to do?

OpenStudy (anonymous):

can you rewrite it using parenthesis? is that x a multiplication sign or a x variable?

OpenStudy (anonymous):

4 and 2/3 times 3 and 5/6

OpenStudy (anonymous):

post a picture and maybe i can solve it

OpenStudy (anonymous):

you want to multiply all those together?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{4}{1} * \frac{2}{3} * \frac{3}{1} * \frac{5}{6}\] now simplify \[\frac{4}{1} * \frac{1}{1} * \frac{1}{1} * \frac{5}{3}\] now multiply all numerator with numerators and denominators with denominators and you get \[\frac{20}{3}\]

OpenStudy (anonymous):

hmm okay i forgot ha ha , sorry thanks so much

OpenStudy (anonymous):

no problems